THe bottom of a 25ft ladder leaning against a house is 7ft from the house. If the ground is level, how high on the house does the ladder reach?

Answer by stanbon(75887)

THe bottom of a 25ft ladder leaning against a house is 7ft from the house. If the ground is level, how high on the house does the ladder reach?
---
Draw the picture.
You have a right triangle with hypotenuse = 25
and base = 7.
---
Use Pythagoras to find the height.
height^2 = hypotenuse^2 - base^2
---
h^2 = 25^2-7^2
h^2 = 576
h = 24 ft.
============
Cheers,
Stan H.